# The Unmet Promise of Synthetic Training Images: Using Retrieved Real Images Performs Better

### v2

@@ Appendix A Broader Impacts
Appendix B Details in Methodology
B.1 Sourcing data by generating synthetic images
@@
Each of these captions are directly used as text input to Stable Diffusion 1.5 to produce our targeted synthetic dataset ${\mathcal{D}_{\mathcal{C}}^{(\text{synthetic})}}$. When sampling from Stable Diffusion, we denoise for 50 DDIM [58] steps starting from Gaussian noise, using a classifier-free guidance [26] scale of 2.0.
@@
Appendix C Details in Experimental Setup
C.1 Retrieval hyperparameters
@@
C.2 Details for downstream benchmarks
@@
C.3 Details for model adaptation
@@
Appendix D Details in Analysis Experiments
D.1 Synthetically perturbing retrieved images
We use SDEdit [42] to synthetically perturb the retrieved real images by adding Gaussian noise and then denoising with Stable Diffusion 1.5. We study the perturbation across 10 different choices of the noise scale (i.e. perturbation strength) parameter $\gamma\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$. All other hyperparameters for the perturbation (e.g. guidance scale, noise scheduler) are fixed to the same hyperparameters used to generate synthetic images from scratch (detailed in Appendix B.1) When using SDEdit to perturb a retrieved real image, we additionaly require a text prompt to guide the denoising. We find that using an empty string for the prompt results in incoherent images. We thus use the string “a photo of {classname}”, where {classname} is the label name assigned to the retrieved image being perturbed. Thus, even at noise scale $\gamma=1.0$ (i.e., beginning the denoising at the endpoint of Stable Diffusion’s forward diffusion process [^2] Even at $\gamma=1.0$, which maps input images to the endpoint of the forward diffusion process, Stable Diffusion’s noise schedule does not convert the input image into pure isotropic Gaussian noise. This is due to the noise schedule parameters.), we would not expect model performance from training on perturbed retrieved images to be identical to model performance from training on our synthetically generated images; our synthetically generated images are synthesized using LLM-written prompts that contain richer information.
Appendix E Details in Licensing Information
E.1 Benchmarks

### v5

@@ Appendix A Broader Impacts
Appendix B Additional Findings
We further validate the main findings of our paper by (1) providing further analysis of the synthetic training data and (2) considering two additional variations upon our experimental setup: applying more aggressive decontamination, and generating synthetic images with alternative prompting strategies.
B.1 Additional data analysis: CLIP score distributions
Do the distributions of CLIP similarity scores in the final synthetic and retrieved training datasets significantly differ? Since we filter the synthetic and retrieved data independently of each other to maintain a clean experimetnal setup (i.e., we do not use any information from the retrieved images to inform the selection of synthetic data, and vice versa), there may be differences in CLIP similarity that potentially explain gaps in downstream training performance.
We histogram the CLIP similarity score distributions of the resulting filtered synthetic and retrieved training data in Figure 8. Overall, despite setting the filter score threshold for synthetic and retrieved data independently, we find that the distribution of post-filtering synthetic image CLIP scores is right-shifted compared to the distribution of post-filtering retrieved image CLIP scores. In other words, synthetic images have comparable or even higher CLIP scores than retrieved images on average; CLIP judges synthetic data to be higher quality on average. CLIP score differences alone do not explain the lagging training performance of synthetic data.
Further corroborating this finding with a case study, we observed that CLIP assigns the synthetic “flute” images in Figure 3 relatively high similarity scores of $0.285$, $0.265$ and $0.263$, despite the fact they they are obviously wrong to the human eye. For reference, a CLIP score of $0.249$ reflects the top $30\%$ of all retrieved ‘flute’ images. Overall, while the current best practice of using CLIP score to filter synthetic data does improve performance, CLIP score remains limited. Future work may explore synthetic data filtering methods that do not depend on CLIP.
Figure 8: We visualize the distribution of CLIP scores in the filtered retrieved and filtered synthetic targeted adaptation datasets. Overall, CLIP assigns synthetic data higher scores on average, despite its lower training performance.
B.2 Additional experiment: further decontaminating retrieved data for benchmark train sets
For our main experiments, we decontaminated all retrieved data for the downstream benchmark test sets to avoid test set leakage. However, retrieving from LAION-2B may also result in images from downstream benchmark training sets being included in the retrieved data. While this contamination will also affect synthetic data, as the generator is trained on the same data we retrieve from, the potential contamination is arguably more explicit when we retrieve directly. We thus further decontaminated all retrieved data for the benchmark train sets following [20]. We report the amount data removed by this additional decontamination step in Table 1 and plot the results of training on twice-decontaminated retrieved data in Figure 9. Overall, while some retrieved images were indeed similar to train set data (an average $1.9\%$ of retrieved data was removed), discarding them minimally impacted performance. Our main findings remain unchanged.
ImageNet-1K DTD FGVC-Aircraft StanfordCars Flowers102
145370 (5.6%) 0 (0.0%) 328 (0.2%) 6647 (3.5%) 342 (0.1%)
Table 1: We report the number of images removed from our original retrieved datasets (which are test set decontaminated) after further decontaminating with respect to the downstream benchmark train set. We report the percentage of the retrieved dataset removed in parentheses.
Figure 9: We further decontaminate our original retrieved data (already test set decontaminated) for the benchmark train set and finetune off-the-shelf CLIP on the train+test decontaminated retrieved data. Overall, train set decontamination has little performance impact; synthetic data remains behind. Note, no train set contamination was found for DTD.
B.3 Additional experiment: generating synthetic data with alternative prompts
In our main experiments, we generate synthetic data via LLM-generated image captions following the SOTA method of SynCLR [62]. SynCLR shows that prompting for synthetic images with LLM captions outperforms many alternative prompting strategies at the hundred-million scale, such as prompting with alt-text from LAION images. To ensure that we utilize the most performant synthetic data generation method available, we sought to validate this finding in our experimental setup. We further compare against the strategy suggested in [37], which involves prompting with BLIP-2 generated captions of real images.
Starting from the unfiltered retrieved training data, we used BLIP-2 to caption each image and construct prompts of the form “a photo of classname, BLIP-2 image caption” following [37]. We then performed top-30% CLIP filtering on the resulting synthetic images. We compare the performance of training with filtered synthetic images generated with three prompting distinct strategies to our filtered retrieved data in Figure 10. Specifically, we compare training with filtered synthetic images generated from (1) our original LLM prompts, (2) BLIP-2 captions, and (3) LAION alt-text from retrieved data. Overall, among the three generation strategies, our original LLM prompts perform best on ImageNet, and perform comparably to BLIP-2 captions and LAION alt-text on Aircraft.
Figure 10: We compare the training performance of targeted synthetic images generated via (1) our original LLM-guided prompts, (2) LAION alt-text from retrieved data, or (3) BLIP-2 captions of retrieved data. Our original LLM prompts yield the best synthetic data for ImageNet; all 3 generation stratgies perform comparably on Aircraft. Regardless of the strategy chosen, synthetic data lags retrieved data.
Appendix C Details in Methodology
C.1 Sourcing data by generating synthetic images
@@
Each of these captions are directly used as text input to Stable Diffusion 1.5 to produce our targeted synthetic dataset ${\mathcal{D}_{\mathcal{C}}^{(\text{synthetic})}}$. When sampling from Stable Diffusion, we denoise for 50 DDIM [60] steps starting from Gaussian noise, using a classifier-free guidance [26] scale of 2.0. We choose Stable Diffusion as our generator because (1) it is pretrained on an open-source dataset, LAION-2B, and (2) to better maintain consistency with recent work [62, 63, 56, 23].
@@
C.2 Filtering synthetic and retrieved data
We rank and filter synthetic data independently, using a per-class threshold to identify and keep the top 30% of images for each class and for each dataset. We adopt CLIP as our synthetic image filter as it is the current best filtering method we are aware of [24].
Appendix D Details in Experimental Setup
D.1 Retrieval hyperparameters
@@
D.2 Details for downstream benchmarks
@@
D.3 Details for model evaluation
When performing zero-shot evaluation, the finetuned model $W\circ\text{CLIP}$ is directly used to inference the downstream test set without any additional tuning. When performing linear probing evaluation, we replace $W$ with a randomly initialized linear head $W^{\prime}$, freeze the CLIP encoder, and train $W^{\prime}$ with a standard cross entropy classification loss on the downstream benchmark train set; we perform inference with $W^{\prime}\circ\text{CLIP}$.
D.4 Details for model adaptation
@@
Appendix E Details in Analysis Experiments
E.1 Synthetically perturbing retrieved images
We use SDEdit [43] to synthetically perturb the retrieved real images by adding Gaussian noise and then denoising with Stable Diffusion 1.5. We study the perturbation across 10 different choices of the noise scale (i.e. perturbation strength) parameter $\gamma\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$. All other hyperparameters for the perturbation (e.g. guidance scale, noise scheduler) are fixed to the same hyperparameters used to generate synthetic images from scratch (detailed in Appendix C.1) When using SDEdit to perturb a retrieved real image, we additionaly require a text prompt to guide the denoising. We find that using an empty string for the prompt results in incoherent images. We thus use the string “a photo of {classname}”, where {classname} is the label name assigned to the retrieved image being perturbed. Thus, even at noise scale $\gamma=1.0$ (i.e., beginning the denoising at the endpoint of Stable Diffusion’s forward diffusion process [^2] Even at $\gamma=1.0$, which maps input images to the endpoint of the forward diffusion process, Stable Diffusion’s noise schedule does not convert the input image into pure isotropic Gaussian noise. This is due to the noise schedule parameters.), we would not expect model performance from training on perturbed retrieved images to be identical to model performance from training on our synthetically generated images; our synthetically generated images are synthesized using LLM-written prompts that contain richer information.
Appendix F Details in Licensing Information
F.1 Benchmarks